OpenStudy (sedatefrog712):

Medal. The lengths of two sides of a triangle are shown below: Side 1: 5x2 + 2x - 3 Side 2: 8x - 3x2 + 2 The perimeter of the triangle is 2x3 - 5x2 + 4x - 8. Part A: What is the total length of the two sides, 1 and 2, of the triangle? Part B: What is the length of the third side of the triangle? Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer.

OpenStudy (akashdeepdeb):

Perimeter of a triangle (P) = A + B + C A = 5x2 + 2x - 3 B = 8x - 3x2 + 2 C = ? P = 2x3 - 5x2 + 4x - 8 See any way to solve this now?

OpenStudy (sedatefrog712):

add like terms?

OpenStudy (sedatefrog712):

@AkashdeepDeb

OpenStudy (sedatefrog712):

@triciaal

OpenStudy (triciaal):

What is the total length of the two sides, 1 and 2, of the triangle? add 1 and 2

OpenStudy (sedatefrog712):

do I add like terms

OpenStudy (triciaal):

Part B: What is the length of the third side of the triangle?

OpenStudy (triciaal):

yes

OpenStudy (akashdeepdeb):

P = A+ B + C C = P - (A + B) Find C.

OpenStudy (sedatefrog712):

-15x^4 + 10x^2 -5 after I add like terms right

OpenStudy (triciaal):

(5x2 + 2x - 3) + (8x - 3x2 + 2) =(5x^2 -3x^2) +(2x + 8x) + (-3 + 2 ) do you see the like terms now?

OpenStudy (sedatefrog712):

yes I see them

OpenStudy (triciaal):

all the x^2 together ; the x terms together etc

OpenStudy (triciaal):

so for Part A write the result when you simplify each group

OpenStudy (sedatefrog712):

-15x^4 + 10x^2 -5 for a

OpenStudy (triciaal):

no you do not have any terms in x^4 review the post x^2 together x together and numbers together

OpenStudy (triciaal):

(5x^2 -3x^2) +(2x + 8x) + (-3 + 2 ) 2 x^2 + 10 x -1

OpenStudy (sedatefrog712):

o ok ok I see

OpenStudy (triciaal):

good we can't mix apples and oranges even though they are both fruits

OpenStudy (sedatefrog712):

True

OpenStudy (triciaal):

now for part b you have the sum of all 3 sides and you know the sum of 2 sides to get the 3rd side do as above P - (a + b) = c where c is the 3rd side and (a + b) is the sum of the (fruit basket)

OpenStudy (triciaal):

2x3 - 5x2 + 4x - 8 - (2 x^2 + 10 x -1) = ?

OpenStudy (sedatefrog712):

o ok so I then distribute it correct

OpenStudy (triciaal):

2 x^3 - 7 x^2 + 14 x -7

OpenStudy (triciaal):

yes and simplify by grouping like terms

OpenStudy (sedatefrog712):

o ok I see

OpenStudy (sedatefrog712):

I will be right back in just a sec

OpenStudy (triciaal):

Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer.

OpenStudy (triciaal):

I am not sure how to determine if the polynomials are closed under addition and subtraction

OpenStudy (sedatefrog712):

its ok I believe that this is closed under addition and subtraction thank you so much for helping me in just a few minutes can you help me with another one but itll be about 15 minuets before I come back is that ok

OpenStudy (triciaal):

post

OpenStudy (sedatefrog712):

ill be back in 15 mins got to eat lunch